# FREAK: Frequency-modulated High-fidelity and Real-time Audio-driven Talking Portrait Synthesis thanks: *Corresponding author

### v1

@@ III-C Loss Functions
$\mathcal{L}_{percep}=\|\mathcal{E}(I_{o})-\mathcal{E}(I_{gt})\|_{2},$ (6)
where $\mathcal{E}$ represents the feature extractor of VGG-19[50].
@@
$\mathcal{L}_{freq}=\|\mathcal{F}(I_{o})-\mathcal{F}(I_{gt})\|_{1},$ (7)
@@
$\mathcal{L}=\lambda_{rec}\mathcal{L}_{rec}+\lambda_{percep}\mathcal{L}_{percep}+\lambda_{freq}\mathcal{L}_{freq},$ (8)
@@ IV Experiments
IV-A Experimental Settings
IV-A1 Dataset
We use data from [6] and [51], selecting 7 videos containing 7 different IDs. The duration of each video ranges from 4 to 6 minutes. The first 80% of each video is used as the training set, and the remaining 20% is reserved for testing.
IV-A2 Comparison Methods
We compare our method with several state-of-the-art approaches including: Wav2Lip[16], DINet[2], MuseTalk[17], IP-LAP[14], RAD-NeRF[1] and SadTalker[3]. Among these, Wav2Lip, MuseTalk and RAD-NeRF are real-time methods. SadTalker belongs to one-shot methods.
IV-A3 Comparison Metrics
The proposed method is evaluated under three scenarios: reconstruction, cross-audio, and one-shot. In the reconstruction scenario, since ground truth of the same identity is available, Frechet Inception Distance (FID) [52] and Frechet Video Distance (FVD) [53] are used for visual quality. Landmark Distance (LMD) [54] is employed to evaluate mouth movement synchronization. In the cross-audio scenario, audio clips from [55], including French and English, are used to drive all the models. As no ground truth of the same identity exists in this case, Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) [16] are adopted to evaluate the synchronization between lip movements and audio without references. The runtime for inferring a single frame serves as the running speed metric.
Figure 3: Qualitative Comparisons with State-of-the-Art Methods. Three identities, each with different speech content, are compared against Wav2Lip, DINet, IP-LAP, MuseTalk, RAD-NeRF and SadTalker. The results of SadTalker are used for comparison with those obtained by our method in the one-shot setting.
IV-A4 Implementation Details
@@
IV-B Quantitative Evaluation
As shown in Table I, our FREAK achieves the best performance in terms of FID, FVD and LMD. The performance on FVD indicates that our results are optimal in terms of both video quality and temporal consistency. Since Wav2Lip is trained with a pre-trained SyncNet[48], it attains exceptionally high scores on both LSE-D and LSE-C, even surpassing the ground truth. However, our method also outperforms most competing approaches in these metrics, demonstrating its strong synchronization capabilities. For LMD, which is calculated by subtracting the mean from each landmark coordinate to obtain a normalized distance, it reflects the accuracy of lip movements to some extent in the one-shot scenario. Compared to SadTalker, a single-image-driven method, our approach also demonstrates significant advantages in FID, FVD and LMD. Notably, SadTalker leverages a pre-trained Wav2Lip model for lip synthesis, resulting in reasonably high scores on the two synchronization metrics.
@@ IV-B Quantitative Evaluation
Figure 4: Results without both VEFM & AVFM in the one-shot setting. The consistency of the head structure between frames is difficult to maintain, with noticeable jitter between frames. Some frames exhibit ghosting or distortion.
Figure 5: The impact of the context length of hubert features on the effect.
IV-C Qualitative Evaluation
@@
TABLE II: Ablation for VEFM, AVFM and Frequency Loss
@@
TABLE III: User study.
@@
IV-D Ablation Study
We conducted ablation experiments on several key components of our model, as shown in Table II. For the variant without VEFM, the features encoded from the reference image and the mask image are directly concatenated, followed by a $1\times 1$ convolution for channel transformation. Similarly, in the absence of AVFM, visual and audio features are directly concatenated. Moreover, we examined the impact of removing the frequency loss, and the effect of changing the context length of the input Hubert features.
@@ IV-D Ablation Study
IV-E User Study
To provide a more comprehensive evaluation of the proposed model, we conduct user studies with 10 participants unrelated to this study on 21 videos generated by ours and the six other methods. We adopt the widely used Mean Opinion Scores (MOS) rating protocol. Each participant is asked to rate from 1-5 for the talking portrait results based on three major aspects: audio-visual synchronization, video quality and realness. Audio-visual sync. focuses on the synchronization between audio and lip movements. Video quality is related to video clarity, facial clarity, and the preservation of textures. Realness is a more comprehensive aspect, indicating the authenticity of the video and determining whether it is AI-generated. The more natural the portrait’s speech performance, the higher the score. We collect the rating results and compute the average score of each method. The statistics are shown in Table III. FREAK surpasses previous methods in all evaluations.
V Discussion and Conclusion

### v2

@@ 3.3. Loss Functions
(6) $\mathcal{L}_{percep}=\|\mathcal{E}(I_{o})-\mathcal{E}(I_{gt})\|_{2},$
where $\mathcal{E}$ represents the feature extractor of VGG-19(Simonyan and Zisserman, 2014).
@@
(7) $\mathcal{L}_{freq}=\|\mathcal{F}(I_{o})-\mathcal{F}(I_{gt})\|_{1},$
@@
(8) $\mathcal{L}=\lambda_{rec}\mathcal{L}_{rec}+\lambda_{percep}\mathcal{L}_{percep}+\lambda_{freq}\mathcal{L}_{freq},$
@@ 4. Experiments
4.1. Experimental Settings
4.1.1. Dataset
We use data from (Guo et al., 2021) and (Zhang et al., 2020), selecting 7 videos containing 7 different IDs. The duration of each video ranges from 4 to 6 minutes. The first 80% of each video is used as the training set, and the remaining 20% is reserved for testing.
4.1.2. Comparison Methods
We compare our method with several state-of-the-art approaches including: Wav2Lip(Prajwal et al., 2020), DINet(Zhang et al., 2023b), MuseTalk(Zhang et al., 2024), IP-LAP(Zhong et al., 2023), RAD-NeRF(Tang et al., 2022) and SadTalker(Zhang et al., 2023a). Among these, Wav2Lip, MuseTalk and RAD-NeRF are real-time methods. SadTalker belongs to one-shot methods.
4.1.3. Comparison Metrics
The proposed method is evaluated under three scenarios: reconstruction, cross-audio, and one-shot. In the reconstruction scenario, since ground truth of the same identity is available, Frechet Inception Distance (FID) (Heusel et al., 2017) and Frechet Video Distance (FVD) (Unterthiner et al., 2019) are used for visual quality. Landmark Distance (LMD) (Chen et al., 2018) is employed to evaluate mouth movement synchronization. In the cross-audio scenario, audio clips from (Shen et al., 2022), including French and English, are used to drive all the models. As no ground truth of the same identity exists in this case, Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) (Prajwal et al., 2020) are adopted to evaluate the synchronization between lip movements and audio without references. The runtime for inferring a single frame serves as the running speed metric.
Figure 3. Qualitative Comparisons with State-of-the-Art Methods. Three identities, each with different speech content, are compared against Wav2Lip, DINet, IP-LAP, MuseTalk, RAD-NeRF and SadTalker. The results of SadTalker are used for comparison with those obtained by our method in the one-shot setting.
4.1.4. Implementation Details
@@
4.2. Quantitative Evaluation
As shown in Table 1, our FREAK achieves the best performance in terms of FID, FVD and LMD. The performance on FVD indicates that our results are optimal in terms of both video quality and temporal consistency. Since Wav2Lip is trained with a pre-trained SyncNet(Chung and Zisserman, 2017), it attains exceptionally high scores on both LSE-D and LSE-C, even surpassing the ground truth. However, our method also outperforms most competing approaches in these metrics, demonstrating its strong synchronization capabilities. For LMD, which is calculated by subtracting the mean from each landmark coordinate to obtain a normalized distance, it reflects the accuracy of lip movements to some extent in the one-shot scenario. Compared to SadTalker, a single-image-driven method, our approach also demonstrates significant advantages in FID, FVD and LMD. Notably, SadTalker leverages a pre-trained Wav2Lip model for lip synthesis, resulting in reasonably high scores on the two synchronization metrics.
@@ 4.2. Quantitative Evaluation
Figure 4. Results without both VEFM & AVFM in the one-shot setting. The consistency of the head structure between frames is difficult to maintain, with noticeable jitter between frames. Some frames exhibit ghosting or distortion.
Figure 5. The impact of the context length of hubert features on the effect.
4.3. Qualitative Evaluation
@@
Table 2. Ablation for VEFM, AVFM and Frequency Loss
@@
Table 3. User study.
@@
4.4. Ablation Study
We conducted ablation experiments on several key components of our model, as shown in Table 2. For the variant without VEFM, the features encoded from the reference image and the mask image are directly concatenated, followed by a $1\times 1$ convolution for channel transformation. Similarly, in the absence of AVFM, visual and audio features are directly concatenated. Moreover, we examined the impact of removing the frequency loss, and the effect of changing the context length of the input Hubert features.
@@ 4.4. Ablation Study
4.5. User Study
To provide a more comprehensive evaluation of the proposed model, we conduct user studies with 10 participants unrelated to this study on 21 videos generated by ours and the six other methods. We adopt the widely used Mean Opinion Scores (MOS) rating protocol. Each participant is asked to rate from 1-5 for the talking portrait results based on three major aspects: audio-visual synchronization, video quality and realness. Audio-visual sync. focuses on the synchronization between audio and lip movements. Video quality is related to video clarity, facial clarity, and the preservation of textures. Realness is a more comprehensive aspect, indicating the authenticity of the video and determining whether it is AI-generated. The more natural the portrait’s speech performance, the higher the score. We collect the rating results and compute the average score of each method. The statistics are shown in Table 3. FREAK surpasses previous methods in all evaluations.
5. Discussion and Conclusion
@@
6. Acknowledgement
This work was supported by the Fundamental Research Funds for the Central Universities of China.